OpenStudy (anonymous):

Find the derivative of f(x) = -2x2 + 11x at x = 9

OpenStudy (trojanpoem):

Derivative of x^n = nx^(n-1) f(x) = -2x^2 + 11x f'(x) = - 4x + 11 at x = 9 (Plug this value) f'(9) = -4 * 9 + 11 = -36 + 11 = -25